A 17-foot piece of string is cut into two pieces so that the longer piece is 2 feet longer than twice the shorter piece. Find the lengths of both pieces.
Answer by checkley77(12844)

x+2x+2=17
3x=17-2
3x=15
x=15/3
x=5 ft. for the short piece.
17-5=12 ft. for the longer piece.
Proof:
5+2*5+2=17
5+10+2=17
17=17